# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

### v1

@@ Abstract
Given a number of datasets for evaluating the performance of single reference methods for the low-lying excited states of closed-shell molecules, a comprehensive dataset for assessing the performance of multireference methods for the low-lying excited states of open-shell systems is still lacking. For this reason, we propose an extension (QUEST#4X) of the radial subset of QUEST#4 [J. Chem. Theory Comput. 2020, 16, 3720] to cover 110 doublet and 39 quartet excited states. Near-exact results obtained by iCIPT2 (iterative configuration interaction with selection and second-order perturbation correction) are taken as benchmark to calibrate SDSCI (static-dynamic-static configuration interaction) and SDSPT2 (static-dynamic-static second-order perturbation theory), which are minimal MRCI and CI-like perturbation theory, respectively. It is found that SDSCI is very close in accuracy to ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles), although its computational cost is just that of one iteration of the latter. Unlike most variants of MRPT2, SDSPT2 treats single and multiple states in the same way, and performs similarly as MS-NEVPT2 (multi-state n-electron valence second-order perturbation theory). These findings put the SDS family of methods (SDSPT2, SDSCI, and iCIPT2, etc.) on a firm basis.
@@ 1 Introduction
The last decades have witnessed fast progresses in the development of both wavefunction- and density-based quantum chemical methods for describing electronic structures of chemical systems. It is generally true that the strength and weakness of each of such methods should be uncovered before it can be applied safely to unknown problems. To this end, some standardized datasets should be established, such that they can be employed to identify the error bars of a given method. A dataset is composed of two ingredients, target systems and reference data. The former refers to chosen molecules and their properties, whereas the latter refers to corresponding experimental or highly accurate theoretical values. The target systems can be classified according to the simple criteria, (1) closed-shell (CS) or open-shell (OS), (2) main group (MG) or transition metal (TM), (3) ground state (GS) or excited state (ES), thereby leading to 8 types of datasets (see Fig. 1), to which the available datasets can be assigned. It can be seen from Table 1 that there exist 123 datasets for ground states of CS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 , OS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 , and CS/OS-TM 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123 (including lanthanides109, 118, 117 and actinides110, 116). Such datasets were mainly used to calibrate SR methods 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139 , especially density functional theory (DFT) 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82 , although some of them were also used to access the performance of MR methods 118, 117, 140, 115, 100, 141, 101, 142, 23, 22. In parallel, there exist 73 datasets oriented to excited states of CS-MG 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 , OS-MG 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 , CS-TM 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 , and OS-TM 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210 , aiming to calibrate both SR 4, 5, 6, 7, 8, 143, 144, 145, 146, 203, 204, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 198, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 199, 200, 201, 202, 205, 206, 207, 208, 209, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 215, 231, 232, 233, 234, 235, 236, 237, 238, 239 and MR 146, 147, 203, 204, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 205, 206, 207, 209, 210, 238, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252 methods.
@@
As for the reference data, the very first point to be realized lies in that experimental measurements often cannot be taken as they stand. For instance, experimentally measured 0-0 transition energies cannot directly be compared to theoretically calculated vertical excitation energies (VEE), which are not observable. Instead, geometric and vibrational effects must be added to theoretical VEEs or subtracted from experimental 0-0 transition energies 253, 143, 171, 153, 156, 179, 166, 254, 176. The lack of sufficient experimental data necessitates the use of some high-level methods as the reference for low-level methods. However, this is somewhat empirical, if not arbitrary. In particular, it is often the case that the chosen high-level method is itself not sufficiently accurate. For instance, the near-exact iCIPT2/TZVP (iterative configuration interaction with selection and second-order perturbation255, 256) calculations147 of the singlet VEEs of the dataset196 overturned the recommendation of using the CASPT2/TZVP rather than the CC3/TZVP values as the theoretical best estimates (TBE),196 which was based simply on the extent of MR characters.
Close inspections of the up-to-date datesets (see Table 1) reveal that there exists only one dataset147 for calibrating MR methods for the excited states of open-shell organic systems. As such, more comprehensive datasets are highly desired. As a first try towards this goal, we start with the QUEST#4 dataset145, which contains 24 organic radicals (see Fig. S1 in the Supporting Information). Since only 51 doublet excited states were reported therein, an immediate extension is to cover more excited states, both doublet and quartet. Such extended dataset, covering 110 doublet and 39 quartet states for the 24 radicals, is to be dubbed QUEST#4X. The VEEs calculated by iCIPT2 will be employed to calibrate MR methods [SS/MS-NEVPT2 (single/muti-state second-order n-electron valence state perturbation theory)257, 258, SDSPT2 (static-dynamic-static (SDS) second-order perturbation theory)259, 260, SDSCI259, and ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles)261], as well as the spin-adapted open-shell TD-DFT262, 263, 264. The SDSCI, SDSPT2, and iCIPT2 methods are first recapitulated in Sec. 2, which is followed by computational details in Sec. 3. The results are analyzed in Sec. 4. The paper is closed with concluding remarks in Sec. 5.
@@ 2 Computational Methods
The (restricted) static-dynamic-static (SDS) family of methods for strongly correlated systems of electrons have been detailed before147, 265. Therefore, it is only necessary to explain here their international relations. In the first place, no matter how many electrons and how many orbitals are to be correlated, SDSCI259 always constructs and diagonalizes a -by- matrix for states simultaneously. Therefore, SDSCI is a minimal MRCI resulting from the combination of the good of perturbation theory, intermediate Hamiltonian, and configuration interaction. Taking the eigenvectors of SDSCI as new references and repeating SDSCI until convergence, we obtain iCI266, where each iteration accesses a space that is higher by two ranks than that of the preceding iteration. That is, up to -tuple excitations (relative to the initial reference space) can be accessed if iterations are carried out. Because of the variational nature, any minor loss of accuracy stemming from the contractions can be removed by carrying out some micro-iterations. In other words, by controlling the numbers of macro- and micro-iterations, iCI will generate a series of contracted/uncontracted single/multireference CISD, with the resulting energy being physically meaningful at each level. It has been shown both theoretically and numerically that iCI can converge monotonically and quickly from above to FCI, even when starting with a very poor initial guess. As such, iCI can be interpreted as an exact solver of FCI. Further combined with the selection of configurations for static correlation and perturbation correction for dynamic correlation, we obtain the near-exact iCIPT2255, 256. On the other hand, if the block of the SDSCI matrix (i.e., matrix elements within the first-order interacting space) is replaced by , we obtain SDSPT2259, 260, a CI-like MRPT2 that treats single and multiple states in the same way and is particularly advantageous when a number of states are nearly degenerate (because of the sufficient relaxation of the reference coefficients)260. In short, if the SDSCI calculation is to be performed, we would obtain SDSPT2 results for free. The latter further yields, e.g., MS-NEVPT2 results for free, because all matrix elements required by MS-NEVPT2258 are already available. Moreover, SDSCI can be taken as a start of ic-MRCISD, so as to facilitate the convergence of the latter. As a matter of fact, the increments of the ic-MRCISD iterations clearly indicate the accuracy of SDSCI. Given so many good features, neither SDSCI nor SDSPT2 is size consistent. However, the errors can readily be cured147 by using the Pople correction267. This is also the case for ic-MRCISD261. Therefore, the SDSPT2, SDSCI, and ic-MRCISD results reported here all refer to those with the Pople corrections.
@@ 3 Computational Details
The 24 radicals and their geometries in the QUEST#4 dataset145 were held unchanged in the iCIPT2255, 256, ic-MRCISD261, SDSCI259, 260, SDSPT2260, and (MS-)NEVPT2257, 258 calculations carried out with the BDF program package268, 269, 270, 271, 272 under the highest Abelian group symmetries. Since the aug-cc-pVTZ (AVTZ) basis sets273, 274, 275 are good enough for most of the excitation energies (with the mean deviation (MD) only of 0.02 eV from the complete basis set (CBS) limit)145, they were also used here. SA-CASSCF (state-averaged complete active space self-consistent field) calculations with equal weights for all states were first carried out (see Sec.4.3 for the various active spaces). The Dyall Hamiltonian276 was then diagonalized within the chosen active space and used as the active part of in both SDSPT2 and (MS-)NEVPT2. The inactive part () of is composed of orbital energies for the doubly and zero occupied orbitals. They were obtained by diagonalizing the generalized Fock matrix (constructed with the state-averaged one-particle density matrix (1RDM)) within the doubly and zero occupied subspaces separately. However, such a choice of is problematic for pure Rydberg states when they were averaged equally with valence states in the SA-CASSCF calculations. The reason is very simple: pure Rydberg states stem from configurations very different from those of valence states. As such, it should be better to use state-specific orbital energies, that is, the doubly and zero occupied orbital energies are to be determined by the Fock matrix that is constructed with the 1RDM of each Rydberg state itself. Conceptually, the resulting state-specific orbitals should also be used when constructing the SDSPT2/MS-NEVPT2 effective Hamiltonian matrix. However, this is not only expensive due to additional integral transformations but also means that the off-diagonal Hamiltonian matrix elements would involve two sets of non-orthogonal orbitals. Therefore, SA-CASSCF instead of state-specific orbitals were still used in such ‘state-dependent Fock’ for 242. Although it is essentially indistinguishable from the ‘sate-averaged Fock’ for valence excitations242, it is a must for pure Rydberg states (see Table S1 in the Supporting Information).
@@
Since iCIPT2 samples the whole Hilbert space, the initial SA-CASSCF calculation is not really needed, although it does provide a good start for the selection procedure. Anyway, iCIPT2 can work with natural orbitals (NOs) obtained by diagonalizing the 1RDM for the selected configurations with, e.g., . Here, is the single parameter involved in iCIPT2. That is, all configuration state functions with coefficients smaller in absolute value than will be pruned away from the variational space determined iteratively (for more details, see Ref. 256). With the fixed NOs, a series of iCIPT2 calculations can be performed with decreasing and are then extrapolated to the zero limit. The extrapolation is particularly robust when a linear fit is sufficient. It turns out that should be used for C3H5, CNO, CON, and CH2NO2, whereas are already sufficient for other systems.
@@
Finally, calculations with the spin-adapted open-shell TD-DFT262, 263, 264 under the Tamm-Dancoff approximation (denoted as X-TDA), in conjunction with the BHandHLYP functional 277, 278, 279 (the simplest yet reliable functional for the low-lying excited states of open-shell systems200, 201), were also performed for comparison.

### v2

@@ Abstract
Given a number of datasets for evaluating the performance of single reference methods for the low-lying excited states of closed-shell molecules, a comprehensive dataset for assessing the performance of multireference methods for the low-lying excited states of open-shell systems is still desired. For this reason, we propose an extension (QUEST#4X) of the radical subset of QUEST#4 [J. Chem. Theory Comput. 2020, 16, 3720] to cover 110 doublet and 39 quartet excited states. Near-exact results obtained by iCIPT2 (iterative configuration interaction with selection and second-order perturbation correction) are taken as benchmark to calibrate SDSCI (static-dynamic-static configuration interaction) and SDSPT2 (static-dynamic-static second-order perturbation theory), which are minimal MRCI and CI-like perturbation theory, respectively. It is found that SDSCI is very close in accuracy to ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles), although its computational cost is just that of one iteration of the latter. Unlike most variants of MRPT2, SDSPT2 treats single and multiple states in the same way, and performs similarly as MS-NEVPT2 (multi-state n-electron valence second-order perturbation theory). These findings put SDSCI and SDSPT2 on a firm basis.
@@ 1 Introduction
The last decades have witnessed fast progresses in the development of both wavefunction- and density-based quantum chemical methods for describing electronic structures of chemical systems. It is generally true that the strengths and weaknesses of each of such methods should be uncovered before they can be applied safely to unknown problems. To this end, some standardized datasets should be established, such that they can be employed to identify the error bars of a given method. A dataset is composed of two ingredients, target systems and reference data. The former refers to chosen molecules and their properties, whereas the latter refers to corresponding experimental or highly accurate theoretical values. The target systems can be classified according to the simple criteria, (1) closed-shell (CS) or open-shell (OS), (2) main group (MG) or transition metal (TM), (3) ground state (GS) or excited state (ES), thereby leading to 8 types of datasets (see Fig. 1), to which the available datasets can be assigned. It can be seen from Table 1 that there exist at least 123 datasets for ground states of CS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 , OS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 , and CS/OS-TM 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123 (including lanthanides109, 118, 117 and actinides110, 116). Such datasets were mainly used to calibrate SR methods 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139 , especially density functional theory (DFT) 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82 , although some of them were also used to assess the performance of MR methods 118, 117, 140, 115, 100, 141, 101, 142, 23, 22. In parallel, there exist 73 datasets oriented to excited states of CS-MG 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 , OS-MG 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 , CS-TM 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 , and OS-TM 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210 , aiming to calibrate both SR 4, 5, 6, 7, 8, 143, 144, 145, 146, 203, 204, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 198, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 199, 200, 201, 202, 205, 206, 207, 208, 209, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 215, 231, 232, 233, 234, 235, 236, 237, 238, 239 and MR 146, 147, 203, 204, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 205, 206, 207, 209, 210, 238, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252 methods.
@@
As for the reference data, the very first point to be realized lies in that experimental measurements often cannot be taken as they stand. For instance, experimentally measured 0-0 transition energies cannot directly be compared to theoretically calculated vertical excitation energies (VEE), which are not observable. Instead, geometric and vibrational effects must be added to theoretical VEEs or subtracted from experimental 0-0 transition energies 253, 143, 171, 153, 156, 179, 166, 254, 176. The lack of sufficient experimental data necessitates the use of some high-level methods as the reference for low-level methods. However, this is not always reliable. It is often the case that the chosen high-level method is itself not sufficiently accurate. For instance, the near-exact iCIPT2/TZVP (iterative configuration interaction with selection and second-order perturbation255, 256) calculations147 of the singlet VEEs of the dataset196 overturned the recommendation of using the CASPT2/TZVP rather than the CC3/TZVP values as the theoretical best estimates (TBE),196 which was based simply on the extent of MR characters.
Close inspections of the up-to-date datasets (see Table 1) reveal that there exists only one dataset147 for calibrating MR methods for the excited states of open-shell organic systems. As such, more comprehensive datasets are highly desired. As a first try towards this goal, we start with the QUEST#4 dataset145, which contains 24 organic radicals (see Fig. S1 in the Supporting Information). Since only 51 doublet excited states were reported therein, an immediate extension is to cover more excited states, both doublet and quartet. Such an extended dataset, covering 110 doublet and 39 quartet states for the 24 radicals, is to be dubbed QUEST#4X. The VEEs calculated by iCIPT2 will be employed to calibrate MR methods [SS/MS-NEVPT2 (single/multi-state second-order n-electron valence state perturbation theory)257, 258, SDSPT2 (static-dynamic-static (SDS) second-order perturbation theory)259, 260, SDSCI259, and ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles)261], as well as the spin-adapted open-shell TD-DFT262, 263, 264. The SDSCI, SDSPT2, and iCIPT2 methods are first recapitulated in Sec. 2, which is followed by computational details in Sec. 3. The results are analyzed in Sec. 4. The paper is closed with concluding remarks in Sec. 5.
@@ 2 Computational Methods
The (restricted) static-dynamic-static (SDS) family of methods for strongly correlated systems of electrons has been presented at length before147, 265. Therefore, it is only necessary to explain here their internal relations. In the first place, no matter how many electrons and how many orbitals are to be correlated, SDSCI259 always constructs and diagonalizes a -by- matrix for states simultaneously. Therefore, SDSCI is a minimal MRCI resulting from the combination of the good of perturbation theory, intermediate Hamiltonian, and configuration interaction. Taking the eigenvectors of SDSCI as new references and repeating SDSCI until convergence, we obtain iCI266, where each iteration accesses a space that is higher by two ranks than that of the preceding iteration. That is, up to -tuple excitations (relative to the initial reference space) can be accessed if iterations are carried out. Because of the variational nature, any minor loss of accuracy stemming from the contractions can be removed by carrying out some micro-iterations. In other words, by controlling the numbers of macro- and micro-iterations, iCI will generate a series of contracted/uncontracted single/multireference CISD, with the resulting energy being physically meaningful at each level. It has been shown both theoretically and numerically that iCI can converge monotonically and quickly from above to FCI, even when starting with a very poor initial guess. As such, iCI can be interpreted as an exact solver of FCI. Further combined with the selection of configurations for static correlation and perturbation correction for dynamic correlation, we obtain the near-exact iCIPT2255, 256. On the other hand, if the block of the SDSCI matrix (i.e., matrix elements within the first-order interacting space) is replaced by , we obtain SDSPT2259, 260, a CI-like MRPT2 that treats single and multiple states in the same way and is particularly advantageous when a number of states are nearly degenerate (because of the sufficient relaxation of the reference coefficients)260. In short, if the SDSCI calculation is to be performed, we would obtain SDSPT2 results for free. The latter further yields, e.g., MS-NEVPT2 results for free, because all matrix elements required by MS-NEVPT2258 are already available. Moreover, SDSCI can be taken as a start of ic-MRCISD, so as to facilitate the convergence of the latter. As a matter of fact, the increments of the ic-MRCISD iterations clearly indicate the accuracy of SDSCI. Given so many good features, neither SDSCI nor SDSPT2 is size consistent. However, the errors can readily be cured147 by using the Pople correction267. This is also the case for ic-MRCISD261. Therefore, the SDSPT2, SDSCI, and ic-MRCISD results reported here all refer to those with the Pople corrections.
@@ 3 Computational Details
The 24 radicals and their geometries in the QUEST#4 dataset145 were held unchanged in the iCIPT2255, 256, ic-MRCISD261, SDSCI259, 260, SDSPT2260, and (MS-)NEVPT2257, 258 calculations carried out with the BDF program package268, 269, 270, 271, 272 under the highest Abelian group symmetries. Since the aug-cc-pVTZ (AVTZ) basis sets273, 274, 275 are good enough for most of the excitation energies (with the mean deviation (MD) only of 0.02 eV from the complete basis set (CBS) limit)145, they were also used here. SA-CASSCF (state-averaged complete active space self-consistent field) calculations with equal weights for all states were first carried out (see Sec.4.3 for the various active spaces). The Dyall Hamiltonian276 was then diagonalized within the chosen active space and used as the active part of the zeroth-order Hamiltonian in both SDSPT2 and (MS-)NEVPT2. The inactive part () of is composed of orbital energies for the doubly and zero occupied orbitals. They were obtained by diagonalizing the generalized Fock matrix (constructed with the state-averaged one-particle density matrix (1RDM)) within the doubly and zero occupied subspaces separately. However, such a choice of is problematic for states dominated by Rydberg characters (more than 70%) when they were averaged equally with valence states in the SA-CASSCF calculations. The reason is very simple: Rydberg states stem from configurations very different from those of valence states. As such, it should be better to use state-specific orbital energies, that is, the doubly and zero occupied orbital energies are to be determined by the Fock matrix that is constructed with the 1RDM of each Rydberg state itself. Conceptually, the resulting state-specific orbitals should also be used when constructing the SDSPT2/MS-NEVPT2 effective Hamiltonian matrix. However, this is not only expensive due to additional integral transformations but also means that the off-diagonal Hamiltonian matrix elements would involve two sets of non-orthogonal orbitals. Therefore, SA-CASSCF instead of state-specific orbitals were still used in such ‘state-dependent Fock’ for 242. Although it is essentially indistinguishable from the ‘state-averaged Fock’ for valence excitations242 , it is a must for pure Rydberg states (see Table S1 in the Supporting Information). Nevertheless, the ‘state-averaged Fock’ still performs better for states of heavily mixed valence and Rydberg characters (e.g., the , , and states of C3H5 are roughly 1:1 mixtures of valence and Rydberg excitations).
@@
Since iCIPT2 samples the whole Hilbert space, the initial SA-CASSCF calculation is not really needed, although it does provide a good start for the selection procedure. Anyway, iCIPT2 can work with natural orbitals (NOs) obtained by diagonalizing the 1RDM for the selected configurations with, e.g., . Here, is the single parameter involved in iCIPT2. That is, all configuration state functions with coefficients smaller in absolute value than will be pruned away from the variational space determined iteratively (for more details, see Ref. 256). With the fixed NOs, a series of iCIPT2 calculations can then be performed with decreasing , so as to approach the FCI limit () by linear fit of the iCIPT2 vs PT2 energy plot. For both the ground and excited states, the extrapolation uncertainties at the 95% confidence level were kept below 1 mHartree. Such uncertainties for the ground and excited states were then summed up as the error bars (less than 2 mHartree or 0.05 eV) for the extrapolated VEEs. The extrapolation distances, i.e., the differences between the largest iCIPT2 calculations (at ) and extrapolated values, are also provided for reference.
@@
Finally, calculations with the spin-adapted open-shell TD-DFT262, 263, 264 under the Tamm-Dancoff approximation (denoted as X-TDA), in conjunction with the BHandHLYP functional 277, 278, 279 (the simplest yet reliable functional for the low-lying doublet-doublet200 and doublet-quartet excitations201 of radicals), were also performed for comparison.